OpenStudy (anonymous):

Put each of the following in vertex form and identify the vertex. y = x^2 + 6x + 3

OpenStudy (anonymous):

y = x^2 + 6x + 3 y= x^2+2(3)x+9 +3-9 y=(x+3)^2-6 y+6=(x+3)^2 vertex is -3,-6